Finding the perpendicular line of Y= -3x -6 with the ordered pair being (-1,5. I need a new equation for the perpendicular line. I also need the slope
Answer by mananth(16946)

1 y = -3 x -6
Divide by 1
y = -3 x + -6
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = -3

The slope of a line perpendicular to the above line will be the negative reciprocal 1/3
Because m1*m2 =-1
The slope of the required line will be 1/3

m= 1/3 ,point ( -1 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = - 1/ 3 + b
b= 16/3
m= 1/3
The required equation is y = (1/ 3) x + 16/ 3